# Order Out of Noise and Disorder: Fate of the Frustrated Manifold

### v1

@@ 4.2 Connection to Instanton Physics in Spin Glasses and QFT Models
Furthermore, the multiplicity of near-degenerate final states in our model (all band orientations on the sphere, all ring configurations on the torus, etc.) is not a discrete set as in simple double-well instantons, but a continuous manifold, the Goldstone manifold associated with the spontaneously broken rotational symmetry. The instanton-like transition thus lands the system somewhere on this manifold, and subsequent dynamics explores it through thermal fluctuations rather than through further instanton transitions. On the sphere, this exploration is purely diffusive ($\alpha\approx 1$); on the torus and cylinder, additional physics (metric structure and boundary conditions) produces anomalous dynamics (superdiffusive and ballistic, respectively).
@@ 4.7 Comparison with Condensed Matter Systems
The isotropic-to-nematic transition in liquid crystals breaks SO(3) $\to$ SO(2), with the director $\mathbf{n}$ selecting a preferred axis, directly analogous to our band normal on the sphere. Both systems exhibit relaxational dynamics toward free energy minima, and both have soft modes with quadratic dispersion characteristic of dissipative systems. The key difference is that nematic ordering is an equilibrium phase transition, while our band formation is a non-equilibrium relaxation to a NESS.
Ferromagnets exhibit spontaneous magnetization that breaks SO(3) $\to$ SO(2), yielding one Type-II magnon with $\omega\propto k^{2}$. The reduction from two broken generators to one Goldstone mode occurs because the nonzero order parameter itself breaks symmetry between transverse directions. Our system exhibits a related structure: once a band forms with a definite orientation, rotations about different axes have inequivalent effects on the configuration.
Experimental systems of colloidal particles on curved interfaces [41] provide a direct physical realization of particles on 2D manifolds. These systems exhibit defect-mediated interactions from Gaussian curvature, frustration between local packing and global topology, and slow relaxation dynamics. Our model differs in having random rather than uniform interactions, but shares the geometric constraints and dissipative dynamics. The dimensional reduction we observe might have experimental signatures in colloidal systems with appropriately engineered interactions.
@@ 4.8 Relation to Self-Organizing Systems
Our model exhibits spontaneous structure formation from initially disordered configurations, inviting comparison with self-organizing systems studied across physics, chemistry, and biology [9, 43, 10]. Classical examples (Bénard convection rolls, Turing patterns, the Belousov-Zhabotinsky oscillator, flocking behavior) typically require: (i) continuous energy or matter flux maintaining non-equilibrium conditions; (ii) nonlinear feedback amplifying fluctuations; (iii) carefully tuned control parameters exceeding critical thresholds; and often (iv) engineered driving protocols.
The structures emerging in our model (equatorial bands, boundary clusters, poloidal rings) are geometrically simpler than the intricate patterns of classical self-organization. Yet they arise through a mechanism that is, in a different sense, more fundamental: no engineered nonlinearities, no fine-tuning of parameters, and no external driving beyond thermal contact with a heat bath. The organizing agents are: (i) quenched disorder in the random couplings $\phi_{ij}$, which creates frustration and a complex energy landscape; (ii) thermal noise, which enables exploration of this landscape and drives instanton-like transitions; and (iii) manifold topology and geometry, which constrain which low-dimensional configurations are kinetically accessible. Structure formation occurs for any non-zero coupling strength and temperature, with no instability threshold to cross.
This robustness suggests that disorder-induced self-organization constitutes a distinct paradigm alongside the instability-driven mechanisms of Prigogine and Haken. The principle “order from noise” articulated by von Foerster [44] acquires a concrete realization: three sources of randomness and constraint (quenched disorder, thermal fluctuations, and geometric topology) interact synergistically to produce ordered structures autonomously. While our emergent configurations are topologically simple (great-circle bands, minor-circle rings), the underlying mechanism is general. More complex combinations of noise statistics, disorder correlations, and manifold topology may lead to richer self-organized structures, still driven by the same autonomous interplay rather than by externally imposed driving or carefully engineered feedback.
@@ 4.9 Synthesis: Unifying Principles
    Zero-torque property enables orientation diffusion: The vanishing of deterministic torque from pairwise interactions means that the orientation of emergent structures (band normal, ring positions) evolves as a pure random walk, explaining the diffusive Goldstone mode dynamics. Inter-structure angular velocity correlations are consistently anti-correlated ($\rho<0$), reflecting conservation of total stochastic angular impulse at each instant.
@@
    Dimensional reduction is generic: Across spin glasses, cosmology, and our model, final configurations have lower effective dimensionality than the starting point, suggesting that frustrated dissipative systems tend to concentrate on submanifolds.
@@
    Spectral signatures reflect local dynamics: The universal $S_{\Theta}(\omega)\propto 1/\omega^{2}$ power spectrum on all three geometries is the fingerprint of underlying Langevin dynamics. However, real-space MSD exponents differ (diffusive on sphere, ballistic on cylinder, superdiffusive on torus), indicating that while the spectrum reflects local stochastic forcing, the MSD is sensitive to global constraints from boundaries and metric structure.
@@
    Geometry controls fluctuation behavior: Structure formation affects rotational fluctuations differently across manifolds: band formation on the sphere constrains fluctuations ($\sigma$ decreases), cluster formation on the cylinder has no effect ($\sigma$ constant), and ring formation on the torus liberates fluctuations ($\sigma$ increases). This pattern reflects how emergent structures interact with manifold geometry to either restrict or expand the accessible phase space.
@@
    Slow dynamics signal landscape complexity: The persistence of aging, metastability, and slow collective modes indicates navigation through rugged energy landscapes, a universal feature connecting spin glasses, structural glasses, and our geometric model.
@@ 5 Conclusion
The title of this paper, echoing Prigogine and Stengers’ celebrated Order out of Chaos [45], encapsulates our central finding: order emerges from the combination of noise and disorder, with the manifold’s topology determining the fate of the frustrated system. As discussed in Section 4, our model differs fundamentally from classical self-organizing systems: it requires no external driving, no fine-tuned parameters, no engineered nonlinearities. The “driving” comes entirely from the thermal bath. Yet from three minimal ingredients (Brownian noise, quenched random couplings, and manifold geometry), ordered structures emerge spontaneously: bands on the sphere, rings on the torus, clusters on the cylinder. Geometry and topology convert randomness into order.
@@
Unlike conventional spontaneous symmetry breaking, the symmetry-breaking direction slowly drifts via thermal noise while the reduced-dimensional structure persists. The slowly evolving collective variables (band orientation, ring positions, cluster locations) constitute classical realizations of dissipative Nambu-Goldstone modes, specifically the type-A diffusion modes predicted for symmetry breaking in open systems [2, 3]. This identification is supported by diffusive mean-square angular displacement ($\alpha\approx 0.9$) and $1/\omega^{2}$ power spectra on the sphere, with the bounded cylinder’s ballistic behavior ($\alpha\approx 2.0$) and the torus’s superdiffusive behavior ($\alpha\approx 1.7$) confirming that diffusive dynamics requires a continuous Goldstone manifold.
@@
The key difference lies in the nature of the ordered state. In equilibrium ferromagnets, the magnetization direction freezes and Goldstone modes are propagating spin waves with dispersion $\omega\propto k^{2}$ (Type-II). In our dissipative system, the collective orientation never freezes but performs a random walk on the Goldstone manifold—the hallmark of Type-A diffusive Goldstone modes in open systems [2, 3]. Both involve macroscopic order emerging from microscopic degrees of freedom through symmetry-breaking transitions, but the fluctuation character (propagating vs. diffusive) reflects whether the system is isolated or coupled to a thermal bath.

### v2

@@ 4.2 Connection to Instanton Physics in Spin Glasses and QFT Models
Furthermore, the multiplicity of near-degenerate final states in our model (all band orientations on the sphere, all ring configurations on the torus, etc.) is not a discrete set as in simple double-well instantons, but a continuous manifold, the Goldstone manifold associated with the spontaneously broken rotational symmetry. The instanton-like transition thus lands the system somewhere on this manifold, and subsequent dynamics explores it through thermal fluctuations rather than through further instanton transitions. Section 3.10 quantifies this exploration: the structure normal $\hat{\mathbf{n}}(t)$ accumulates roughly $2300^{\circ}$ of precession on the sphere (at rate $\sim 61^{\circ}$/time), $2400^{\circ}$ on the torus ($\sim 65^{\circ}$/time), and $1200^{\circ}$ on the cylinder ($\sim 32^{\circ}$/time) during the NESS phase. The correlation time $\tau_{c}\approx 18$ on the sphere and torus sets the timescale for orientation decorrelation, far exceeding the microscopic thermal correlation time $\tau_{v}<0.2$.
@@ 4.7 Comparison with Condensed Matter Systems
The isotropic-to-nematic transition in liquid crystals breaks SO(3) $\to$ SO(2), with the director $\mathbf{n}$ selecting a preferred axis, directly analogous to our band normal on the sphere. Both systems exhibit relaxational dynamics toward free energy minima, and both have soft modes with quadratic dispersion characteristic of dissipative systems. Our structure orientation analysis (Section 3.10) makes this analogy precise: the band normal $\hat{\mathbf{n}}(t)$ plays the role of the nematic director, exploring the Goldstone manifold $S^{2}\cong\mathrm{SO}(3)/\mathrm{SO}(2)$ with correlation time $\tau_{c}\approx 18$ and accumulating roughly $2300^{\circ}$ of precession over the NESS phase. The key difference is that nematic ordering is an equilibrium phase transition with a frozen director below $T_{c}$, while our band normal never freezes but continues to diffuse indefinitely, driven by thermal noise in the NESS.
Ferromagnets exhibit spontaneous magnetization that breaks SO(3) $\to$ SO(2), yielding one Type-II magnon with $\omega\propto k^{2}$. The reduction from two broken generators to one Goldstone mode occurs because the nonzero magnetization itself breaks symmetry between transverse directions: fluctuations of $\mathbf{M}$ perpendicular to $\langle\mathbf{M}\rangle$ are coupled through the commutation relations $M_{x},M_{y}]\propto M_{z}$. Our system exhibits a related structure. Once a band forms with normal $\hat{\mathbf{n}}$, rotations about different axes have inequivalent effects on the configuration. The measured isotropic orientation statistics ($\sigma(n_{x})\approx\sigma(n_{y})\approx\sigma(n_{z})\approx 0.1$$0.3$ on the sphere) suggest that our non-equilibrium steady state explores all directions equivalently over long times, in contrast to the frozen magnetization direction in a ferromagnet. The power spectrum $S_{n_{z}}(\omega)\propto\omega^{-1.6}$ on the sphere is shallower than the $1/\omega^{2}$ expected for type-A diffusion modes, possibly reflecting the bounded nature of the orientation variable $\hat{\mathbf{n}}$ on $S^{2}$.
Experimental systems of colloidal particles on curved interfaces [45] provide a direct physical realization of particles on 2D manifolds. These systems exhibit defect-mediated interactions from Gaussian curvature, frustration between local packing and global topology, and slow relaxation dynamics. Our model differs in having random rather than uniform interactions, but shares the geometric constraints and dissipative dynamics. The dimensional reduction we observe, quantified by the structure thickness $\mathcal{T}$ decreasing from $\sim 0.6$ to $\sim 0.3$ as the band forms (Table 13), might have experimental signatures in colloidal systems with appropriately engineered interactions.
@@ 4.8 Relation to Self-Organizing Systems
Our model exhibits spontaneous structure formation from initially disordered configurations, inviting comparison with self-organizing systems studied across physics, chemistry, and biology. The field encompasses several distinct paradigms with different mechanisms and requirements.
Dissipative structures.
The program initiated by Prigogine [12] and developed by Haken [47] and others [13] focuses on pattern formation in systems maintained far from equilibrium by continuous energy or matter flux. Classical examples (Bénard convection rolls, Turing patterns, the Belousov-Zhabotinsky oscillator) require: (i) sustained thermodynamic gradients; (ii) nonlinear feedback that amplifies fluctuations; and (iii) control parameters tuned above critical thresholds. These “dissipative structures” are intrinsically non-equilibrium phenomena that vanish when the external driving ceases.
Self-organized criticality.
Bak and collaborators [48, 49] introduced a different paradigm: systems that spontaneously evolve toward a critical state without fine-tuning of parameters. The canonical example is a sandpile that, through slow addition of grains, organizes itself to a critical slope where avalanches of all sizes occur with power-law statistics. SOC systems share with our model the absence of parameter fine-tuning, but differ in requiring slow external input (grain addition) and producing scale-free, intermittent dynamics (avalanches) rather than persistent geometric structures.
Our model: disorder-induced organization.
The structures emerging in our model (equatorial bands, boundary clusters, poloidal rings) arise through a mechanism distinct from both paradigms above. Our model requires: no sustained thermodynamic gradients (only thermal equilibrium with a heat bath); no slow external input (the quenched couplings are fixed, not continuously supplied); and no parameter fine-tuning (structure formation occurs for any non-zero coupling strength and temperature). The organizing agents are: (i) quenched disorder in the random couplings $\phi_{ij}$, which creates frustration and a complex energy landscape; (ii) thermal noise, which enables exploration of this landscape and drives instanton-like transitions; and (iii) manifold topology and geometry, which constrain which low-dimensional configurations are kinetically accessible.
This combination suggests that disorder-induced self-organization constitutes a distinct paradigm. The principle “order from noise” articulated by von Foerster [50] acquires a concrete realization: quenched disorder, thermal fluctuations, and geometric topology interact to produce ordered structures autonomously. Unlike dissipative structures, our patterns persist in thermal equilibrium (the NESS is maintained by the heat bath alone). Unlike SOC, our system produces specific geometric configurations determined by topology rather than scale-free avalanche statistics. The emergent order is geometrically simple (great-circle bands, minor-circle rings) but arises from minimal ingredients without external orchestration.
@@ 4.9 Synthesis: Unifying Principles
    Zero-torque property enables orientation diffusion: The vanishing of deterministic torque from pairwise interactions means that the orientation of emergent structures evolves as a pure random walk. The structure normal $\hat{\mathbf{n}}(t)$ diffuses on its Goldstone manifold with geometry-dependent correlation times: $\tau_{c}\approx 18$ on the sphere and torus (continuous structures), but $\tau_{c}\approx 3$ on the cylinder (discrete clusters). This order-of-magnitude difference reflects the greater susceptibility of discrete configurations to thermal reorientation.
@@
    Dimensional reduction is generic: Across spin glasses, cosmology, and our model, final configurations have lower effective dimensionality than the starting point, suggesting that frustrated dissipative systems tend to concentrate on submanifolds. The structure thickness $\mathcal{T}$ quantifies this reduction: on the sphere, $\mathcal{T}$ decreases from $0.62$ (diffuse) to $0.29$ (well-defined band) during formation.
@@
    Spectral signatures are geometry-dependent: Both MSD exponents and power spectrum exponents vary across geometries. The structure normal power spectrum $S_{n_{z}}(\omega)\propto\omega^{-\beta}$ shows $\beta\approx 1.6$ on the sphere, $\beta\approx 2.0$ on the cylinder, and $\beta\approx 1.4$ on the torus. Only the cylinder approaches the $1/\omega^{2}$ scaling predicted for type-A diffusion modes. The MSD exponents similarly vary (diffusive on sphere, ballistic on cylinder, superdiffusive on torus). This geometry dependence indicates that both local stochastic forcing and global constraints from boundaries, metric structure, and the bounded nature of $\hat{\mathbf{n}}$ on $S^{2}$ shape collective dynamics.
@@
    Geometry controls orientation statistics: The structure normal exhibits geometry-dependent behavior in the NESS. On the sphere, $\hat{\mathbf{n}}$ explores all directions isotropically ($\langle n_{z}\rangle\approx 0$, $\sigma(n_{z})\approx 0.11$). On the cylinder, $\hat{\mathbf{n}}$ is confined to the $xy$-plane ($\langle n_{z}\rangle\approx 0$, $\sigma(n_{z})\approx 0.07$). On the torus, $\hat{\mathbf{n}}$ shows a strong $z$-bias with large fluctuations ($\langle n_{z}\rangle\approx 0.73$, $\sigma(n_{z})\approx 0.44$). These statistics encode how manifold geometry constrains the accessible configurations of the symmetry-breaking order parameter.
@@
    Slow dynamics signal landscape complexity: The persistence of aging, metastability, and slow collective modes indicates navigation through rugged energy landscapes, a universal feature connecting spin glasses, structural glasses, and our geometric model. The long correlation times $\tau_{c}\approx 18$ on the sphere and torus (compared to thermal velocity correlation times $\tau_{v}<0.2$) confirm the adiabatic separation between fast microscopic dynamics and slow collective evolution.
@@ 5 Conclusion
The title of this paper, echoing Prigogine and Stengers’ celebrated Order out of Chaos [51], encapsulates our central finding: order emerges from the combination of noise and disorder, with the manifold’s topology determining the fate of the frustrated system. As discussed in Section 4, our model represents a distinct paradigm of self-organization. Unlike dissipative structures, it requires no sustained thermodynamic gradients. Unlike self-organized criticality, it requires no slow external input and produces specific geometric configurations rather than scale-free avalanches. The only “driving” is thermal contact with a heat bath, and structure formation occurs without parameter fine-tuning. From three minimal ingredients (Brownian noise, quenched random couplings, and manifold geometry), ordered structures emerge spontaneously: bands on the sphere, rings on the torus, clusters on the cylinder. Geometry and topology convert randomness into order.
@@
Unlike conventional spontaneous symmetry breaking, the symmetry-breaking direction slowly drifts via thermal noise while the reduced-dimensional structure persists. The slowly evolving collective variables (band orientation, ring positions, cluster locations) constitute classical realizations of dissipative Nambu-Goldstone modes, specifically the type-A diffusion modes predicted for symmetry breaking in open systems [5, 6]. This identification is supported by diffusive mean-square angular displacement ($\alpha\approx 0.9$) on the sphere. However, the power spectrum exponents are geometry-dependent: $S_{n_{z}}(\omega)\propto\omega^{-\beta}$ with $\beta\approx 1.6$ on the sphere, $\beta\approx 2.0$ on the cylinder, and $\beta\approx 1.4$ on the torus. Only the cylinder approaches the theoretical $1/\omega^{2}$ prediction for type-A diffusion modes. The bounded cylinder’s ballistic MSD ($\alpha\approx 2.0$) and the torus’s superdiffusive behavior ($\alpha\approx 1.7$) confirm that diffusive dynamics requires a continuous Goldstone manifold.
@@
The key difference lies in the nature of the ordered state. In equilibrium ferromagnets, the magnetization direction freezes and Goldstone modes are propagating spin waves with dispersion $\omega\propto k^{2}$ (Type-II). In our dissipative system, the structure normal never freezes but performs a random walk on the Goldstone manifold with correlation time $\tau_{c}\approx 18$—the hallmark of Type-A diffusive Goldstone modes in open systems [5, 6]. Both involve macroscopic order emerging from microscopic degrees of freedom through symmetry-breaking transitions, but the fluctuation character (propagating vs. diffusive) reflects whether the system is isolated or coupled to a thermal bath.